My class has been hitting simplifying down to y=mx+b through equation. For example, 2y+6x=2. However I have hit this question: 19+2x=2y. The teacher has not taught us how to do this and there are no textbooks to reference. Can you send me the answer to 19+2x=2y along with the way to solve it?
Answer by Alan3354(69443)

Can you send me the answer to 19+2x=2y along with the way to solve it?
------------------
You solve for y.
19+2x=2y
2y = 2x + 19
Divide by 2
y = x + 19/2